# SN 2021foa: Deriving a continuity between SN IIn and SN Ibn

### v1

@@ 5 Spectral and Lightcurve modelling
Figure 9 shows the zoomed-in spectral evolution of the line profiles of H, and He i 5876, 6678, 7065 Å. H shows a very complex profile throughout the evolution. Initially, H shows a narrow line on the top of a broad component. Around 3.2 d, we see a narrow P-Cygni component appearing on top of a broad H profile. Thereafter, the H profile is complex and highly asymmetric. Post 14.2 d, the red part of H starts developing, possibly due to contamination from He i 6678 Å. The He i 5876 Å line also shows evolution, with the FWHM varying between 2500 – 4000 km s-1. The narrow notch on top of the He i profile is mostly due to Na I D. The He i 5876 Å grows in strength and by 7.2 d its luminosity becomes comparable to that of H. The He i 7065 Å line develops later, at 14.2 d, and grows in strength thereafter. Interestingly, the FWHM of the He i 5876 Å line is similar to that of the H line throughout the evolution, which again may indicate a mixed composition of the CSM. The implications of these line profiles with regards to the geometry of the SN is discussed in Sect. 5.1.
@@
To discern the origin of the narrow P-Cygni profile of H, we selected the broad emission and absorption components as a continuum and normalize the spectra with respect to it. Figure 10 shows that in the continuum normalized spectra, from 10.8 d to 49.5 d, we still see a narrow emission of FWHM 500 km s-1. No absorption features are seen in the spectra before 8.3 d. The narrow P-Cygni absorption starts to appear at 5.8 d with its blue edge reaching up to velocities of 1000 km s-1. The narrow P-Cygni becomes really prominent at 7.2 d, with the blue edge extending up to 2000 km s-1. The FWHM of the P-Cygni profile typically varied between 500 km s-1 to 800 km s-1.
@@
To compare the origin of the appearance of narrow P-cygni absorption/emission profiles in interacting SNe, we compared the time evolution of the appearance of the P-Cygni profiles for our group of SNe IIn and SNe Ibn (see Figure 11).
@@
For most of the events in our comparison sample (SNe 1996al, 2005la, 2010al, 2016jbu, 2019uo, 2019wep), the narrow P-Cygni lines is present from the beginning of the observed evolution of the SN. This arises due to the presence of pre-shock CSM present along the line of sight (Smith, 2017). However, some objects (SNe 2006jc, 2010mc, 2010al, 2011hw) show only narrow emission profiles. In this case, we are not seeing the ionized pre-shock CSM along the line of sight. In addition to that, SN 2015bh shows a delayed onset of a P-Cygni profile similar to what is seen for SN 2021foa (Elias-Rosa et al., 2016; Thöne et al., 2017), and SN 2009ip shows narrow emission early on, followed by narrow P-Cygni at intermediate times and then narrow emission lines again at a later stage of the evolution. Smith et al. (2009) suggested that in SN 2009ip, narrow P-Cygni initially arose when the star was in the LBV phase. Just after the explosion, it showed narrow emission lines due to interaction with a dense CSM and thereafter from interacting with another shell moving at 2000 km s-1. Smith et al. (2009) also associated the origin of the narrow P-Cygni as due to outbursts a few decades prior to a “hyper eruption" or the final core-collapse. For the case of SN 2021foa, we also see that the velocity of the blue edge extends up to 2000 km s-1 and we see narrow P-Cygni features developing, which indicate the presence of a shell/disk of CSM along the line of sight. A detailed interpretation of this associated with the geometry of the CSM is also described in Sect. 7.1.
@@ 5.1 H decomposition
From the typical explosion circumstances of interacting SNe (Smith, 2017), we expect a narrow component from the unshocked CSM, an intermediate width component either from the e-scattering of the narrow line photons or from the cold, dense shell (CDS), a broad width component from the uninterrupted ejecta which sometimes may show an associated absorption component as well. Therefore, we try to deconstruct H profile in terms of these components. The H profile was fitted with combinations of different line profiles in order to reproduce the overall line profile seen in SN 2021foa at different stages of its evolution. We used i) A narrow Gaussian component that is slightly redshifted from the center ii) A Lorentzian/Gaussian intermediate width component that is redshifted from the center, iii) A broad Gaussian component in absorption. At early times, the overall H profile is better represented by a Lorentzian intermediate width component as the emission lines are dominated by electron scattering. On top of that, we have a narrow emission mostly arising from pre-shocked CSM. During the middle phases, the narrow emission is replaced by a narrow P-Cygni profile, and at late times (30 d), a Gaussian intermediate width component better reproduces the overall line profile as it evolves into a more complex multi-component structure. The choice of the continuum is very critical for performing these fits. The continuum is selected to be far from the line region by at least 50 Å. The continuum is varied between 50 few Å to check the consistency of the fits. The spectral evolution and the corresponding fits at representative epochs are shown in Figure 12. The parameters of the components obtained from the fitting at all epochs are given in Table 1. The errors listed represent only fitting errors, and other uncertainties like resolution matching, subtraction of host-spectra, and imperfections in wavelength calibration have not been included.
@@ 5.2 Radius and Temperature Evolution
As the SN ejecta is expanding, the shock breakout from the surface of the progenitor is followed by a rapid cooling due to the rapid expansion driven by the shock (Falk et al., 1977). This would lead to a rapidly increasing photospheric radius and a decrease in the temperature of the SN ejecta within a couple of hours of the explosion. However, this will be affected for extended stars with larger radius. For the case of interacting SNe, this is not always the case as the ejecta is masked by CSM (Irani et al., 2023). Figure 14 shows the radius and temperature evolution of SN 2021foa. SN 2021foa first shows a decrease in temperature evolution from 14000 K to about 8000 K. We then see a slight rise in the temperature evolution of SN 2021foa from 7300 K to 8300 K between 17 d and 23 d post maximum, during the shoulder in the lightcurve, and then the temperature evolution becomes flatter. This rise in the temperature indicates an injection of energy in the cooling ejecta, perhaps due to interaction with additional CSM or interaction with regions of enhanced CSM density. This is in turn affected by opacity effects in the CSM ejecta interacting zone.
The black body radius of SN 2021foa increases from 5500 R⊙ to 15500 R⊙ at 10 d past maximum light. From 7 d to 20 d past max, the radius stays at 16000 R⊙. The radius evolution then shows a small shoulder similar to the temperature evolution, fluctuating between the values of 16000 R⊙ to 15000 R⊙ and then decreases to a value of 2800 R⊙. The late time radius evolution is flat, as is the temperature evolution and the luminosity as well.
@@ 5.3 Hydrodynamical Modelling
We conducted light-curve modeling of SN 2021foa using the one-dimensional multi-frequency radiation hydrodynamics code STELLA (Blinnikov et al., 1998, 2000, 2006). Because STELLA treats radiation hydrodynamics in multi-frequencies, STELLA can construct pseudo-bolometric light curves that can be directly compared with the observed ones.
@@
We found that the CSM with two power-law density components can reproduce the overall light-curve properties of SN 2021foa. The wind velocity is assumed to be . The inner CSM component has . The CSM with can account for the early-time luminosity around the light-curve peak. After around 30 days, the luminosity decline becomes faster than expected from interaction with a CSM with . We found that the fast luminosity decline can be reproduced when the CSM density structure follows from . This CSM component has mass of .
@@
Also, we want to mention that the model is stable around the peak time, however, the latter part of the lightcurve is not well-established due to numerical limitations of STELLA. Hence, a two zone CSM for SN 2021foa well-reproduces the lightcurve until 80 d.
@@ 6 Mass-loss rates
The mass-loss rates are governed by the ejecta-CSM interaction in SNe IIn/Ibn and can be estimated from spectral profiles as well (2020MNRAS.499..129G). Assuming that the luminosity of the ejecta-CSM interaction is fed by energy at the shock front, the progenitor mass-loss rate can be calculated using the relation of Chugai & Danziger (1994):
@@
Since we see narrow emission lines of both H and He i at different stages of the evolution, we assume a typical unshocked wind velocity as observed for LBV winds of 100 km s-1 and 1000 km s-1 for the WR stars. The shock velocity is inferred from the intermediate width component. We want to remark, however, that the first phase might be affected by electron scattering, and there may be contamination by the ejecta signatures at these epochs. We take the shock velocities to be 3612 km sec-1 and 7843 km sec-1 at 5.8 d and 7.2 d. Using the bolometric luminosity at day 5.8 (L erg sec-1), wind speeds of 100 (and 1000) km sec-1 and assuming 50 conversion efficiency (), the estimated mass-loss rate is found to be 0.05 (and 0.5) M⊙ yr-1. The estimated mass-loss rate at 7.2 d when the narrow P-Cygni line profile becomes prominent in the spectra are estimated similarly for the wind speed of 100 (and 1000) km sec-1 and assuming L=5.48 1042 erg sec-1 gives 0.007 (and 0.07) M⊙ yr-1.
@@
In between the forward shock and the reverse shock, there is a contact discontinuity between the shocked CSM and shocked ejecta where material cools, mixes via Rayleigh-Taylor instabilities, and piles up. This is often called the cold, dense shell (CDS), and we define the velocity at this front as vCDS. If we look at the narrow component of the H evolution, it did not change significantly over the evolution of the SN however, VCDS changed at this boundary of 7.2 d in the H evolution. Considering these two epochs of time at 5.8 d and 7.2 d, estimating the time periods preceding the explosion when the progenitor showed some activity is calculated by t = tobs x (vw/vCDS). Assuming vw from the narrow component of H at these two phases indicates that the progenitor had undergone a change in the eruptive activity at two stages corresponding to 0.5 – 1 year before exploding as an SN. This is much lower than the traditional SNe IIn and super luminous SNe IIn like 2006gy, 2010jl and 2017hcc (Smith & McCray, 2007; Jencson et al., 2016; Smith & Andrews, 2020) which showed eruptive activity 6-12 years before the explosion. On the contrary, this is quite similar to SN 2019uo (Gangopadhyay et al., 2020), resulting in a shorter-lived light curve as SNe Ibn. This has also been observed previously for the case of SN 2020oi which is a SN Ic showing radio interaction signatures and showed progenitor activity/ expelled shells 1 yr before the explosion Maeda et al. (2021).
7 Discussion:
@@
In this paper, we present the unique case of SN 2021foa, where we see line luminosity ratios intermediate between SNe IIn and SNe Ibn. We also come to the conclusion that it is H line luminosity that better separates the two populations, while the two classes can not be segregated based on He i 5876 Å line luminosities. SN 2021foa exhibits the classic evolution of line profile shapes that is common in strongly interacting SNe IIn, which transition from symmetric Lorentzian profiles at early times (before and during peak), to irregular, broader, and asymmetric shapes at late times well after peak. The phenomena is understood as a shift from narrow CSM lines broadened by electron scattering to emission lines formed in the post-shock cold dense shell CDS (Smith, 2017; Dessart et al., 2015). In addition, for SN 2021foa, we also see asymmetric He i lines which broaden over time from 2000 km sec-1 to 5000 km sec-1 and show line luminosities comparable to H around the lightcurve peak.
@@
The spectral evolution also shows narrow and intermediate-width H lines at pre-maximum times. Around 3.8 d, we see narrow P-Cygni appearing in H. The narrow P-Cygni in H appears later than the narrow P-Cygni in H. Additionally, a recent study by Ishii et al. (2024) explores the relations between the line shapes and CSM structure by Monte Carlo radiative transfer codes. They find that a narrow line exhibits a P-Cygni profile only when an eruptive mass-loss event forms the CSM. The CSM structure from a steady mass loss will have a negative velocity gradient after the SN event due to radiative acceleration. Therefore, an H photon emitted at the deeper CSM layers, traveling outwards, will never be able to undergo another H transition. However, if there is an eruptive mass-loss that comes into play after a steady wind scenario, then a positive velocity gradient would give rise to narrow P-Cygni lines formed along the line of sight. We see from subsection 5.3 and Section 6, the mass-loss rates typically varies from 10-3 - 10-1 M⊙ yr-1 during this phase. Also, the light curve modelling helps us to infer that the density-radius variation occurs from r-2 to r-5. This validates our scenario that probably the change in this mass-loss rates and possible eruption would have given rise to this P-Cygni arising along the line of sight. The changing mass-loss rates and the eruptive activity seen at 0.5 – 5 years before the explosion typically also govern the dual peaked light curve and appearance of P-Cygni seen in the light curve and spectral evolution of SN 2021foa. Reguitti et al. (2024) checked the pre-cursor activities of SN 2021foa and have detected it only upto 50 days pre-explosion. This implies that even though there is an activity in the progenitor driving the peak luminosities at 0 and 17 d post maximum, it was not significant enough to be detected as a precursor in the light curve a year before the explosion. Elias-Rosa et al. (2016); Thöne et al. (2017) have explained the origin of double-peak due to the interaction with a shell at a later point in time, which can also be a possible case of SN 2021foa. But, we find that a disk-like geometry better explains our overall line profiles, and we discuss below in detail the physical scenario (see subsection 7.1) governing the CSM and the explosion geometry.
7.1 Physical scenario and asymmetry:

### v3

@@ 5 Spectral and Lightcurve modelling
Figure 7 shows the zoomed-in spectral evolution of the line profiles of H, and He i 5876, 6678, 7065 Å. H shows a very complex profile throughout the evolution. Initially, H shows a narrow line on the top of a broad component. Around 3.2 d, we see a narrow P-Cygni component appearing on top of a broad H profile. Thereafter, the H profile is complex and highly asymmetric. Post 14.2 d, the red part of H starts developing, possibly due to contamination from He i 6678 Å. In the inset plot in second panel, we see a narrow He i 5876 Å component followed by the dip which is most likely due to Na I D. The He i 5876 Å line shows evolution, with the FWHM varying between 2500 – 4000 km s-1 which is the narrow to intermediate width component in accordance with Pastorello et al. (2008b). The He i 5876 Å grows in strength and by 7.2 d its luminosity becomes comparable to that of H. The He i 7065 Å line develops later, at 14.2 d, and grows in strength thereafter. Interestingly, the FWHM of the He i 5876 Å line is similar to that of the H line throughout the evolution, which again may indicate a mixed composition of the CSM. The implications of these line profiles with regards to the geometry of the SN is discussed in Sect. 5.1.
@@
To discern the origin of the narrow P-Cygni profile of H, we selected the broad emission and absorption components as a continuum and normalize the spectra with respect to it. Figure 8 shows that in the continuum normalized spectra, from 10.8 d to 49.5 d, we still see a narrow emission of FWHM 500 km s-1. We want to remark that N II lines as well as the narrow H and H lines appears to be redshifted by nearly 100 km sec-1 in all the the high-resolution spectra, including our spectra from SUBARU. Since it is noticed in multiple spectra from different telescopes, it is likely not an issue with the wavelength calibration and is most likely associated with the incorrect estimation of the redshift of the SN/host-galaxy. However, we kept the redshift same with (Reguitti et al., 2024; Farias, 2024) to be consistent. No absorption features are seen in the spectra before 8.3 d. The narrow P-Cygni absorption starts to appear at 5.8 d with its blue edge reaching up to velocities of 1000 km s-1. The narrow P-Cygni becomes really prominent at 7.2 d. The FWHM of the P-Cygni profile typically varied between 500 km s-1 to 800 km s-1.
@@
To compare the origin of the appearance of narrow P-cygni absorption/emission profiles in interacting SNe, we compared the time evolution of the appearance of P-Cygni profiles for our group of SNe IIn and SNe Ibn (see Figure 9).
@@
For most of the events in our comparison sample (SNe 1996al, 2005la, 2010al, 2016jbu, 2019uo, 2019wep), the narrow P-Cygni lines is present from the beginning of the observed evolution of the SN. This arises due to the presence of pre-shock CSM along the line of sight (Smith, 2017). However, some objects (SNe 2006jc, 2010mc, 2010al, 2011hw) show only narrow emission profiles. In this case, we are not seeing the ionized pre-shock CSM along the line of sight. In addition to that, SN 2015bh shows a delayed onset of a P-Cygni profile similar to what is seen for SN 2021foa (Elias-Rosa et al., 2016; Thöne et al., 2017), and SN 2009ip shows narrow emission early on, followed by narrow P-Cygni at intermediate times and then narrow emission lines again at a later stage of the evolution. Smith et al. (2009) suggested that in SN 2009ip, narrow P-Cygni initially arose when the star was in the LBV phase. Just after the explosion, it showed narrow emission lines due to interaction with a dense CSM and thereafter from interacting with another shell moving at 2000 km s-1. Smith et al. (2009) also associated the origin of the narrow P-Cygni as due to outbursts a few decades prior to a “hyper eruption" or the final core-collapse. For the case of SN 2021foa, we also see that the velocity of the blue edge extends up to 1000 km s-1 and we see narrow P-Cygni features developing, which indicate the presence of a shell/disk of CSM along the line of sight. A detailed interpretation of this associated with the geometry of the CSM is also described in Sect. 7.1.
@@ 5.1 H decomposition
From the typical explosion circumstances of interacting SNe (Smith, 2017), we expect a narrow component from the unshocked CSM, an intermediate width component either from the e-scattering of the narrow line photons or from the cold, dense shell (CDS), a broad width component from the uninterrupted ejecta which sometimes may show an associated absorption component as well. Therefore, we try to deconstruct the H profile in terms of these components. The H profile was fitted with combinations of different line profiles in order to reproduce the overall line profile seen in SN 2021foa at different stages of its evolution. We used i) A narrow Gaussian component that is slightly redshifted from the center ii) A Lorentzian/Gaussian intermediate width component that is redshifted from the center, iii) A broad Gaussian component in absorption. At early times, the overall H profile is better represented by a Lorentzian intermediate width component as the emission lines are dominated by electron scattering. On top of that, we have a narrow emission component of H mostly arising from pre-shocked CSM. During the middle phases, the narrow emission is replaced by a narrow P-Cygni profile, and at late times (30 d), a Gaussian intermediate width component better reproduces the overall line profile as it evolves into a more complex multi-component structure. The choice of the continuum is very critical for performing these fits. The continuum is selected to be far from the line region by at least 50 Å. The continuum is varied between 50 few Å to check the consistency of the fits. The spectral evolution and the corresponding fits at representative epochs are shown in Figure 10. The parameters of the components obtained from the fitting at all epochs are given in Table 1. The errors listed represent only fitting errors, and other uncertainties like resolution matching, subtraction of host-spectra, and imperfections in wavelength calibration have not been included.
@@ 5.2 Radius and Temperature Evolution
As the SN ejecta is expanding, the shock breakout from the surface of the progenitor is followed by a rapid cooling due to the rapid expansion driven by the shock (Falk et al., 1977). This would lead to a rapidly increasing photospheric radius and a decrease in the temperature of the SN ejecta within a couple of hours of the explosion. However, this will be affected for extended stars with larger radius. For the case of interacting SNe, this is not always the case as the ejecta is masked by CSM (Irani et al., 2023). Figure 12 shows the radius and temperature evolution of SN 2021foa. The temperature and radius are obtained from the blackbody fits including the UV to IR data. For estimating the values, Stefan-Boltzman law was used to estimate the corresponding parameters. SN 2021foa first shows a decrease in temperature evolution from 14000 K to about 8000 K. We then see a slight rise in the temperature evolution of SN 2021foa from 7300 K to 8300 K between 17 d and 23 d post maximum, during the shoulder in the lightcurve, and then the temperature evolution becomes flatter. This rise in the temperature indicates an injection of energy in the cooling ejecta, perhaps due to interaction with additional CSM or interaction with regions of enhanced CSM density. This is in turn affected by opacity effects in the CSM ejecta interacting zone.
The black body radius of SN 2021foa increases from 5500 R⊙ to 15500 R⊙ at 10 d past maximum light. From 7 d to 20 d past max, the radius stays at 16000 R⊙. The radius evolution then shows a small shoulder similar to the temperature evolution, fluctuating between the values of 16000 R⊙ to 15000 R⊙ and then decreases to a value of 2800 R⊙. The late time radius evolution is flat, as is the temperature evolution and the luminosity as well, but we do not show that in the plot because blackbody approximation fails there.
@@ 5.3 Hydrodynamical Modelling
To construct the bolometric light curve of SN 2021foa between UV to IR bands, we used the SuperBol code (Nicholl, 2018). The missing UV and NIR data was supplemented by extrapolating the Specrtral Energy Distributions (SEDs) using the blackbody approximation and direct integration method as described in Lusk & Baron (2017). A linear extrapolation was performed in the UV regime at late times. We conducted light-curve modeling of SN 2021foa using the one-dimensional multi-frequency radiation hydrodynamics code STELLA (Blinnikov et al., 1998, 2000, 2006). Because STELLA treats radiation hydrodynamics in multi-frequencies, STELLA can construct pseudo-bolometric light curves that can be directly compared with the observed ones.
@@
We found that the CSM with two power-law density components can reproduce the overall light-curve properties of SN 2021foa. The wind velocity is assumed to be adopted from the blue edge of the narrow absorption component of the CSM (see Figure 9). The inner CSM component has . The CSM with can account for the early-time luminosity around the light-curve peak. After around 30 days, the luminosity decline becomes faster than expected from interaction with a CSM with . We found that the fast luminosity decline can be reproduced when the CSM density structure follows from . This CSM component has mass of .
@@
Also, we want to mention that the model is stable around the peak time, however, the latter part of the lightcurve is not well-established due to numerical limitations of STELLA. Hence, a two zone CSM for SN 2021foa well-reproduces the lightcurve until 80 d. We want to remark that even if we obtain a disk-like CSM geometry from spectral modelling in Section 5, our simplistic approximations of spherical geometry well reproduces the order of magnitude of the obtained physical parameters in lightcurve modelling.
@@ 6 Mass-loss rates
The mass-loss rates are governed by the ejecta-CSM interaction in SNe IIn/Ibn and can be estimated from spectral profiles as well (Gangopadhyay et al., 2020). Assuming that the luminosity of the ejecta-CSM interaction is fed by energy at the shock front, the progenitor mass-loss rate can be calculated using the relation of Chugai & Danziger (1994):
@@
Since we see narrow emission lines of both H and He i at different stages of the evolution, we assume a typical unshocked wind velocity as observed for LBV winds of 100 km s-1 and 1000 km s-1 for the WR stars. The shock velocity is inferred from the IW component. We want to remark, however, that the first phase might be affected by electron scattering, and there may be contamination by the ejecta signatures at these epochs. We take the shock velocities to be 3612 km sec-1 and 7843 km sec-1 at 5.8 d and 7.2 d. Using the bolometric luminosity at day 5.8 (L erg sec-1), wind speeds of 100 (and 1000) km sec-1 and assuming 50 conversion efficiency (), the estimated mass-loss rate is found to be 0.05 (and 0.5) M⊙ yr-1. The estimated mass-loss rate at 7.2 d when the narrow P-Cygni line profile becomes prominent in the spectra are estimated similarly for the wind speed of 100 (and 1000) km sec-1 and assuming L=5.48 1042 erg sec-1 gives 0.007 (and 0.07) M⊙ yr-1.
@@
In the region between the forward shock and the reverse shock, there exists a contact discontinuity. This boundary separates the shocked CSM and the shocked ejecta. At this interface, material cools, mixes via Rayleigh-Taylor instabilities, and accumulates. This region is referred to as the cold, dense shell (CDS). The velocity at this boundary is denoted as vCDS. When examining the narrow component of the H line, its evolution did not show significant changes over the lifetime of the SN. However, vCDS showed noticeable variation at the boundary around 7.2 days in the H evolution.
Considering these two epochs of time at 5.8 d and 7.2 d, estimating the time periods preceding the explosion when the progenitor showed some activity is calculated by t = tobs x (vw/vCDS). Assuming vw from the narrow component of H at these two phases indicates that the progenitor had undergone a change in the eruptive activity at two stages corresponding to 0.5 – 1 year before exploding as an SN. This is much lower than the traditional SNe IIn and super luminous SNe IIn like 2006gy, 2010jl and 2017hcc (Smith & McCray, 2007; Jencson et al., 2016; Smith & Andrews, 2020) which showed eruptive activity 6-12 years before the explosion. On the contrary, this is quite similar to SN 2019uo (Gangopadhyay et al., 2020), resulting in a shorter-lived light curve as SNe Ibn. This has also been observed previously for the case of SN 2020oi which is a SN Ic showing radio interaction signatures and showed progenitor activity/ expelled shells 1 yr before the explosion Maeda et al. (2021).
7 Discussion
@@
In this paper, we present the unique case of SN 2021foa, where we see line luminosity ratios intermediate between SNe IIn and SNe Ibn. We also come to the conclusion that it is the H line luminosity that better separates the two populations, while the two classes cannot be segregated based on He i 5876 Å line luminosities. SN 2021foa exhibits the classic evolution of line profile shapes that is common in strongly interacting SNe IIn, which transition from symmetric Lorentzian profiles at early times (before and during peak), to irregular, broader, and asymmetric shapes at late times well after peak. The phenomena is understood as a shift from narrow CSM lines broadened by electron scattering to emission lines formed in the post-shock cold dense shell CDS (Smith, 2017; Dessart et al., 2015). In addition, for SN 2021foa, we also see asymmetric He i lines which broaden over time from 2000 km sec-1 to 5000 km sec-1 and show line luminosities comparable to H around the lightcurve peak.
@@
The spectral evolution also shows narrow and intermediate-width H lines at pre-maximum times. Around 3.8 d, we see narrow P-Cygni appearing in H. The narrow P-Cygni in H appears later than the narrow P-Cygni in H. Furthermore, a recent study by Ishii et al. (2024) explores the relations between line shapes and CSM structure using Monte Carlo radiative transfer codes. They find that a narrow line exhibits a P-Cygni profile only when an eruptive mass-loss event forms the CSM. The CSM structure from a steady mass loss will have a negative velocity gradient after the SN event due to radiative acceleration. Therefore, an H photon emitted at the deeper CSM layers, traveling outwards, will never be able to undergo another H transition. However, if there is an eruptive mass-loss that comes into play after a steady wind scenario, then a positive velocity gradient would give rise to narrow P-Cygni lines formed along the line of sight. We see from subsection 5.3 and Section 6, the mass-loss rates typically varies from 10-3 - 10-1 M⊙ yr-1 during this phase. Also, the light curve modelling helps us to infer that the density-radius variation occurs from r-2 to r-5. This validates our scenario that probably the change in this mass-loss rates and possible eruption would have given rise to this P-Cygni arising along the line of sight. The changing mass-loss rates and the eruptive activity seen at 0.5 – 5 years before the explosion typically also govern the dual peaked light curve and appearance of P-Cygni seen in the light curve and spectral evolution of SN 2021foa. Reguitti et al. (2024) checked the pre-cursor activities of SN 2021foa and have detected it only upto 50 days pre-explosion. This implies that even though there is an activity in the progenitor driving the peak luminosities at 0 and 17 d post maximum, it was not significant enough to be detected as a precursor in the light curve a year before the explosion. Elias-Rosa et al. (2016); Thöne et al. (2017) have explained the origin of double-peak due to the interaction with a shell at a later point in time, which can also be a possible case of SN 2021foa. But, we find that a disk-like geometry better explains our overall line profiles, and we discuss below in detail the physical scenario (see subsection 7.1) governing the CSM and the explosion geometry.
7.1 Physical scenario and asymmetry